help me out

Velocity v = גf, wavelength ג = v/f     and Relative frequency = frequency difference

Notes

C

D

E

F

G

A

B

C

Frequency

264

297

330

352

396

440

495

528

Rel. Freq

?

?

?

?

?

?

?

Wavelength

?

?

?

?

?

?

?

?

?

?

?

?

?

?

?
Answer #1

relative frequency is just the difference of frequency

297-264=33

330-297=33

352-330=22

396-352=44

440-396=44

495-440=55

528-495=33

Wavelength =V/f

f=> calculated above for related frequency

for queries do ask in comments